In general brass is made of two parts of copper to one part zinc. How many pounds of zinc are needes to produce 45 pounds of brass.

Answer by stanbon(75887)

In general brass is made of two parts of copper to one part zinc. How many pounds of zinc are needes to produce 45 pounds of brass.
-----
Note: one to two is the same as x to 2x where x is amt. of zinc; 2x is
amt of copper.
---------------------
Equation:
x + 2x = 45
3x = 45
x = 15 (amt. of zinc needed)
2x = 30 (amt. of copper needed)
===============
Cheers,
Stan H.